OpenStudy (anonymous):

what is the simplified form of x^2/y^3 / x^7/y^6

OpenStudy (anonymous):

\[\huge \cfrac{\cfrac{x^2}{y^3}}{\cfrac{x^7}{y^6}}=\frac{x^2y^6}{x^7y^3}\] now use this rule: \[\LARGE \frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (anonymous):

i honestly dont understand

OpenStudy (anonymous):

part one or part two?

OpenStudy (anonymous):

\[\LARGE \frac{x^2y^6}{x^7y^3}=x^{2-7}y^{6-3}=x^{-5}y^3=\frac{1}{x^5}\cdot y^3\] \[\LARGE =\frac{y^3}{x^5}\]

OpenStudy (anonymous):

where did the one come from?

OpenStudy (anonymous):

draw your question. Let's see what was the given ! ;)